40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Apr 7 11:47:22 2022, Yehonathan, Update, BHD, Initial BHD modeling: AS - LO mode matching 
    Reply  Thu Apr 7 21:15:04 2022, Yehonathan, Update, BHD, Initial BHD modeling: AS - LO mode matching 
       Reply  Mon May 16 16:13:01 2022, Yehonathan, Update, BHD, Initial BHD modeling: Damped suspension model Damped_SUS.pngETMY_SUSPOS_GOL.pdfSUS_Response.png
          Reply  Thu Jun 16 17:40:15 2022, Yehonathan, Update, BHD, Comparison of MICH OLTF model with measurement MICH_Model_Measurement_Comparison.pdf
             Reply  Fri Jun 17 12:05:32 2022, Yehonathan, Update, BHD, Comparison of MICH OLTF model with measurement 
                Reply  Fri Jun 17 13:07:08 2022, Koji, Update, BHD, Comparison of MICH OLTF model with measurement 
Message ID: 16761     Entry time: Thu Apr 7 11:47:22 2022     Reply to this: 16766
Author: Yehonathan 
Type: Update 
Category: BHD 
Subject: Initial BHD modeling: AS - LO mode matching 

I begin modeling the initial BHD setup using Finesse. I started with copying C1_w_BHD.kat from the 40m/bhd repo and making changes to reflect the current BHD setup:

1. OMCs were removed.

2. Only 1 PD per BHD port was left.

3. Transmission of PR2 was changed to 2.2%. The PRG was calculated to be ~15.5.

4. Actual RoCs of new optics were dialed in (Yesterday me and Paco went into the cleanroom to measure the RoCs and they seem to match the datasheets).

Here's a table comparing the old (design?) RoCs with the new RoCs:

  New RoC Old RoC
LO1 5m 6m
LO2 inf inf
LO3 500mm 750mm
LO4 150mm -450mm
AS1 2m 2.8m
AS2 inf inf
AS3 200mm -2m
AS4 750mm 600mm
PR2 2000m -700m
PR3 1000m 1000m
SR3 1000m -700m

 

The changes looked quite alarming, especially for LO4 and AS3, so I wrote a script to calculate the mode matching between the LO and AS beams called AS_LO_ModeMatching.ipynb and pushed it into the repo. In the notebook a bright AS beam is created by creating a small asymmetry between the arms of ~ 0.003 degrees (~10pm). Amplitude detectors were put at the input ports of the BHD BS to calculate the fields in the AS and LO beams. In particular TEM00, TEM02 and TEM20 were measured for each beam.

The calculation shows that with the old RoCs the mode matching between the LO and AS beams is 99% while for the new RoCs it is ~ 50%.

ELOG V3.1.3-